# Synergizing Motion and Appearance: Multi-Scale Compensatory Codebooks for Talking Head Video Generation

### v1

@@ Abstract
Talking head video generation aims to generate a realistic talking head video that preserves the person’s identity from a source image and the motion from a driving video. Despite the promising progress made in the field, it remains a challenging and critical problem to generate videos with accurate poses and fine-grained facial details simultaneously. Essentially, facial motion is often highly complex to model precisely, and the one-shot source face image cannot provide sufficient appearance guidance during generation due to dynamic pose changes. To tackle the problem, we propose to jointly learn motion and appearance codebooks and perform multi-scale codebook compensation to effectively refine both the facial motion conditions and appearance features for talking face image decoding. Specifically, the designed multi-scale motion and appearance codebooks are learned simultaneously in a unified framework to store representative global facial motion flow and appearance patterns. Then, we present a novel multi-scale motion and appearance compensation module, which utilizes a transformer-based codebook retrieval strategy to query complementary information from the two codebooks for joint motion and appearance compensation. The entire process produces motion flows of greater flexibility and appearance features with fewer distortions across different scales, resulting in a high-quality talking head video generation framework. Extensive experiments on various benchmarks validate the effectiveness of our approach and demonstrate superior generation results from both qualitative and quantitative perspectives when compared to state-of-the-art competitors.  [^1] https://shaelynz.github.io/synergize-motion-appearance/
@@ 1 Introduction
Given a source image and a driving video, talking head video generation (Hong et al., 2022; Tao et al., 2024) aims to animate the person in the source image using the pose and expression from the driving video. Due to its widespread applications, such as video conferencing, the film industry, and virtual reality, it has attracted growing interest in the community.
Significant progress has been made on this task in terms of both quality and robustness in recent years. Existing works primarily focus on learning more accurate motion estimation and appearance representation in 2D or 3D to enhance generation quality. Along the direction, unsupervised methods target predicting local motion flows around unsupervised keypoints without relying on facial priors (Siarohin et al., 2019b; Zhao & Zhang, 2022; Wang et al., 2024), and methods based on predefined models (e.g., 3DMM) (Zakharov et al., 2019; Zhang et al., 2023; Ha et al., 2020) focus on learning robust decoding features to generate high-quality face outputs. Despite the promising achievements, critical challenges persist: 1) Some motion patterns (e.g., local subtle motions) cannot be inferred from a single image pair solely relying on unsupervised keypoints or predefined models for motion estimation, as such models often have limited power of motion representation and may fail to capture certain dynamic aspects of the facial motion from single image pairs (see Fig. 1b). 2) Even with accurate motion estimation, highly dynamic and complex motions in driving videos can create ambiguity during generation, as a still source image lacks sufficient appearance information to handle occluded regions or subtle expression changes (see Fig. 1d). This results in noticeable artifacts and a significant drop in the quality of the generated output. Therefore, generating realistic-looking facial images requires not only inferring accurate motion flow between the two given facial images but also compensating for the intermediate appearance decoding features from the one-shot source image for the final generation of face images.
@@
In this work, we aim to synergize motion and appearance by simultaneously learning accurate motion flows for facial warping and robust facial appearance features for face image decoding, to advance talking head generation. Inspired by the success of codebook learning (Van Den Oord et al., 2017) where compact and useful representations of certain modalities are learned from the whole training dataset, we propose to use such representations as additional knowledge and compensate for motion and appearance with them. Specifically, we propose a unified framework that can achieve joint learning of both motion and appearance codebooks with multi-scale compensation. To refine the motion flow between two facial images (i.e., source and driving), we design a multi-scale motion codebook that captures diverse motion patterns across scales from the entire dataset during training. To enhance intermediate warped facial feature maps for image decoding, we introduce a multi-scale appearance codebook that represents diverse appearance patterns learned from the entire dataset. To use the motion and appearance codebooks, we further introduce a transformer-based compensation structure to iteratively refine motion flows in a coarse to fine manner and refine the warped features across different scales. This approach enables us to improve motion accuracy and capture more facial details by leveraging the diverse motion and appearance information stored in the codebooks. To enhance the learning and compensation of both codebooks, we propose a joint training strategy in which the motion and appearance codebooks are learned simultaneously with the entire framework. This approach allows both codebooks to be optimized together, utilizing gradients from the refined warped features to strengthen their mutual influence and improve overall performance. By learning both multi-scale motion and appearance codebooks, our framework refines the motion flow to accurately warp the source facial features, which are further compensated with additional details from the appearance codebook. This process yields robust intermediate facial decoding features, resulting in improved generation.
@@
  • We propose a novel framework that jointly learns multi-scale motion and appearance codebooks. The motion codebook captures motion patterns at varying levels of granularity, while the appearance codebook stores representative facial structure and texture features. This joint learning enables the model to effectively compensate both motion and appearance for advanced generation.
@@
  • Extensive experiments including on challenging datasets demonstrate that our method not only effectively compensates for facial motions and appearances but also significantly outperforms state-of-the-art approaches, generating more realistic talking head videos.
@@ 2 Related Work
Talking Head Video Generation. Existing works on talking head video generation generally separate the motion estimation module and image generation module to disentangle appearance and motion. To transfer the motion, some works require facial priors provided by a pre-trained model during generation. For example, landmark-based approaches (Ha et al., 2020; Zhao et al., 2021; Hong et al., 2024; 2023; Wu et al., 2018; Zakharov et al., 2020) detect pre-defined facial landmarks to transfer the facial pose and expression from a driving frame to the source image. Some other methods (Ren et al., 2021; Zeng et al., 2023; Yao et al., 2020) use parameters from 3D face models (Blanz & Vetter, 2023; Feng et al., 2021; Zhu et al., 2017) as motion descriptors to disentangle identity and pose. However, they normally cannot describe non-facial parts such as hair and neck, and their generation quality is limited by the pre-trained model performance. To address the issue, several methods that do not require any prior knowledge from pre-trained models are proposed. Monkey-Net (Siarohin et al., 2019a) learns sparse motion-related keypoints in an unsupervised manner to describe object movements. FOMM (Siarohin et al., 2019b) extends it with local affine transformation assumption around the keypoints to model complex motion. Subsequent works introduce more flexible mathematical models such as thin-plane spline transformation (Zhao & Zhang, 2022) and continuous piecewise-affine-based transformation (Wang et al., 2024) to increase motion estimation accuracy. Despite the expressiveness of the motion models, they cannot fully describe large head poses and delicate expression changes. MRFA (Tao et al., 2024) tackles the problem by building a correlation volume for each image pair and using it to refine the coarse motion flow iteratively. However, it only uses the warped image feature and a plain image generator for image generation, which may fail when facing extreme pose change as the appearance information from the one-shot source image is not enough. Some works (Yin et al., 2022; Oorloff & Yacoob, 2023; Bounareli et al., 2023) leverage the remarkable generative power of pre-trained StyleGAN (Karras et al., 2019; 2020) for better image generation, but they usually have to balance the editability and fidelity. MCNet (Hong & Xu, 2023) learns a meta-memory bank of spatial facial features to compensate for the warped source features. Different from previous works, we compensate for both motion flows and warped source features with jointly learned multi-scale motion and appearance codebooks to boost the generation quality.
@@
Codebook Learning. Codebook learning aims to learn useful discrete representations with a fixed size. The learned codebook contains rich and compact information, which can facilitate various tasks such as image classification (Cai et al., 2010; Zhang et al., 2009), image synthesis (Esser et al., 2021; Wang et al., 2023b; Chang et al., 2022), blind face restoration (Zhou et al., 2022; Gu et al., 2022) and audio-driven talking head video generation (Wang et al., 2023a). Traditional methods often learn a codebook with unsupervised clustering such as k-means (Csurka et al., 2004). VQ-VAE (Van Den Oord et al., 2017) first incorporates vector quantization in a Variational Autoencoder to learn a codebook containing discrete representative input features. To build a context-rich codebook for images, VQGAN (Esser et al., 2021) further increases the compression rate and adds a discriminator and a perceptual loss on images reconstructed with the codes from the codebook. A transformer is later used to model the composition of the codes for high-resolution image synthesis. CodeFormer (Zhou et al., 2022) uses a learned codebook of compressed high-quality face image features as discrete prior and predicts the code sequence based on the low-quality facial input for blind face restoration. LipFormer (Wang et al., 2023a) learns two codebooks of the upper half face and the bottom half face respectively and predicts the lip codes from the input audio to generate a face video from the audio. We also adopt the idea of codebook learning, but we simultaneously learn multi-scale motion and appearance codebooks that store diverse motion and appearance patterns from the entire dataset during training to facilitate high-quality talking head video generation. The codebooks and the entire framework are trained together so that patterns useful for talking head video generation can be stored in and retrieved from the codebooks.
@@ 3 The Proposed Method
In this section, we will present the details of our framework. We jointly learn multi-scale motion and appearance codebooks with compensation for the motion and intermediate appearance features during generation. We adopt the Taylor expansion approximation method to learn the initial motion flow and the warping manner as same as Siarohin et al. (2019a) for video generation.
@@ 3.2.1 Multi-scale Code Allocation
We aim to refine the initial motion flow and the warped source features across all $N$ scales with multi-scale motion and appearance codebooks. As the scale increases, larger appearance features require finer motion flows for accurate warping and more detailed appearance information for appearance compensation. To provide sufficient motion and appearance compensation at larger scales, we introduce a code allocation scheme for multi-scale motion and appearance codebooks, which divides the codebook into multiple groups and allocates more codes for larger scales. We illustrate the scheme with the motion codebook in Fig. 4.
The motion codebook $\mathcal{C}_{M}=\{\mathbf{m}_{k}\in\mathbb{R}^{d_{m}}\}_{k=1}^{K}$ contains $K$ codes, and we perform motion compensation at $N$ different scales. The $K$ codes are split into $N$ equal groups, each with $K/N$ codes. At scale $i$, the first $i$ groups, totaling $N_{m}^{i}=i\times K/N$ codes, are allocated for motion compensation. Similarly, for the appearance codebook $\mathcal{C}_{A}=\{\mathbf{a}_{k}\in\mathbb{R}^{d_{a}}\}_{k=1}^{T}$ containing $T$ codes, we also allocate the first $i$ groups, which are the first $N_{a}^{i}=i\times T/N$ codes for appearance compensation. This allows codes with smaller indices to capture general motion patterns or coarse appearance patterns shared across scales, while codes with larger indices to focus on finer motion patterns or facial details needed for larger scales. This maximizes the use of the two codebooks by sharing general information across scales while reserving space for scale-specific details.
@@ 3.2.2 Joint Codebook Learning
Motion Codebook Learning. At scale $i$, we use a CNN-based motion flow encoder $\mathcal{E}_{M}$ to map the input motion flow $\mathbf{M}^{i-1}\in\mathbb{R}^{h\times w\times 2}$ into a compact motion flow feature $\mathbf{F}_{m}^{i-1}\in\mathbb{R}^{h_{m}\times w_{m}\times d_{m}}$ where each unit is of length $d_{m}$ and captures a local motion flow pattern on $\mathbf{M}^{i-1}$. We quantize each of its spatial element $\mathbf{F}_{m}^{i-1}(x,y)$ with its nearest code from $\mathcal{C}_{M}^{i}$ to obtain a quantized motion flow feature $\hat{\mathbf{F}}_{m}^{i-1}$:
@@
$\begin{split}\mathcal{L}_{vq,m}^{i}=\lambda_{recon,m}||\hat{\mathbf{M}}^{i-1}-sg[\mathbf{M}^{i-1}]||_{1}&+||sg[E_{M}({\mathbf{M}}^{i-1})]-\hat{\mathbf{F}}_{m}^{i-1}||_{2}^{2}\\ &+\beta||sg[\hat{\mathbf{F}}_{m}^{i-1}]-E_{M}(sg[\mathbf{M}^{i-1}])||_{2}^{2},\end{split}$ (3)
where $sg[\cdot]$ denotes the stop gradient operator, and $\lambda_{recon,m}$ and $\beta$ are the loss term weights. The first term represents the motion flow reconstruction loss, while the last two terms form a code-level loss (Van Den Oord et al., 2017) that minimizes the distance between the latent motion flow units and the motion codes. We stop the gradient of $\mathbf{M}^{i-1}$ to ensure that motion codebook learning at scale $i$ does not interfere with the training of the motion flow estimator or the compensated motion flows from prior scales. The overall loss function for motion codebook learning across all $N$ scales is $\mathcal{L}_{vq,m}=\sum_{i=1}^{N}\mathcal{L}_{vq,m}^{i}$.
@@ 3.2.3 Multi-scale Joint Codebook Compensation
Motion and Appearance Codebook Compensation Coupling. With the multi-scale motion and appearance codebooks, we couple motion and appearance codebook compensation at each scale to produce better image decoding features, as shown in Fig. 2. Specifically, at each scale, motion codebook compensation (MCC) refines the input motion flow $\mathbf{M}^{i-1}$ to obtain the compensated motion flow $\mathbf{M}^{i}$, which is used to warp the source feature $\mathbf{F}_{s}^{i}$. Appearance codebook compensation (ACC) then refines the warped source feature to produce the compensated appearance feature $\mathbf{F}_{c}^{i}$. If $i<N$, $\mathbf{M}^{i}$ is used as the input motion flow for the next scale. In this way, motion and appearance codebook compensation are coupled across scales, enabling more consistent motion and appearance for high-quality video generation.

### v2

@@ Abstract
Talking head video generation aims to generate a realistic talking head video that preserves the person’s identity from a source image and the motion from a driving video. Despite the promising progress made in the field, it remains a challenging and critical problem to generate videos with accurate poses and fine-grained facial details simultaneously. Essentially, facial motion is often highly complex to model precisely, and the one-shot source face image cannot provide sufficient appearance guidance during generation due to dynamic pose changes. To tackle the problem, we propose to jointly learn motion and appearance codebooks and perform multi-scale codebook compensation to effectively refine both the facial motion conditions and appearance features for talking face image decoding. Specifically, the designed multi-scale motion and appearance codebooks are learned simultaneously in a unified framework to store representative global facial motion flow and appearance patterns. Then, we present a novel multi-scale motion and appearance compensation module, which utilizes a transformer-based codebook retrieval strategy to query complementary information from the two codebooks for joint motion and appearance compensation. The entire process produces motion flows of greater flexibility and appearance features with fewer distortions across different scales, resulting in a high-quality talking head video generation framework. Extensive experiments on various benchmarks validate the effectiveness of our approach and demonstrate superior generation results from both qualitative and quantitative perspectives when compared to state-of-the-art competitors. The project page is available at https://shaelynz.github.io/synergize-motion-appearance/.
@@ 1 Introduction
Given a source image and a driving video, talking head video generation [16, 25] aims to animate the person in the source image using the pose and expression from the driving video. Due to its widespread applications, such as video conferencing, the film industry, and virtual reality, it has attracted growing interest in the community.
Recent years have witnessed significant advancements in quality and robustness for this task. Current approaches primarily focus on improving motion estimation accuracy and appearance representation, whether in 2D or 3D, to enhance generation quality. Along the direction, unsupervised methods target predicting local motion flows around unsupervised keypoints without relying on facial priors [24, 43, 27], and methods based on predefined models (e.g., 3DMM) [37, 40, 12] focus on learning robust decoding features to generate high-quality face outputs. Despite the promising achievements, critical challenges persist: 1) Some motion patterns (e.g., local subtle motions) cannot be inferred from a single image pair solely relying on unsupervised keypoints or predefined models for motion estimation, as such models often have limited power of motion representation and may fail to capture certain dynamic aspects of the facial motion from single image pairs (see Fig. 1b). 2) Even with accurate motion estimation, highly dynamic and complex motions in driving videos can create ambiguity during generation, as a still source image lacks sufficient appearance information to handle occluded regions or subtle expression changes (see Fig. 1d). This results in noticeable artifacts and a significant drop in the quality of the generated output. Therefore, generating realistic-looking facial images requires not only inferring accurate motion flow between the two given facial images but also compensating for the intermediate appearance decoding features from the one-shot source image for the final generation of face images.
@@
In this work, we aim to synergize motion and appearance by simultaneously learning accurate motion flows for facial warping and robust facial appearance features for face image decoding, to advance talking head generation. Inspired by the success of codebook learning [26] where compact and useful representations of certain modalities are learned from the whole training dataset, we propose to use such representations as additional knowledge and compensate for motion and appearance with them. Specifically, we propose a unified framework that can achieve joint learning of both motion and appearance codebooks with multi-scale compensation. To refine the motion flow between two facial images (i.e., source and driving), we design a multi-scale motion codebook that captures diverse motion patterns across scales from the entire dataset during training. To enhance intermediate warped facial feature maps for image decoding, we introduce a multi-scale appearance codebook that represents diverse appearance patterns learned from the entire dataset. To use the motion and appearance codebooks, we further introduce a transformer-based compensation structure to iteratively refine motion flows in a coarse to fine manner and refine the warped features across different scales. This approach enables us to improve motion accuracy and capture more facial details by leveraging the diverse motion and appearance information stored in the codebooks. To enhance the learning and compensation of both codebooks, we propose a joint training strategy in which the motion and appearance codebooks are learned simultaneously with the entire framework. This approach allows both codebooks to be optimized together, utilizing gradients from the refined warped features to strengthen their mutual influence and improve overall performance. By learning both multi-scale motion and appearance codebooks, our framework refines the motion flow to accurately warp the source facial features, which are further compensated with additional details from the appearance codebook. This process yields robust intermediate facial decoding features, resulting in improved generation.
@@
  • We propose a novel framework that jointly learns multi-scale motion and appearance codebooks. The motion codebook captures motion patterns at varying levels of granularity, while the appearance codebook stores representative facial structure and texture features. This joint learning enables the model to effectively compensate for both motion and appearance for advanced generation.
@@
  • Extensive experiments, including those on challenging datasets, demonstrate that our method not only effectively compensates for facial motions and appearances but also significantly outperforms state-of-the-art approaches, generating more realistic talking head videos.
@@ 2 Related Work
Talking Head Video Generation. Existing works on talking head video generation generally separate the motion estimation module and the image generation module to disentangle appearance and motion. To transfer the motion, some works require facial priors provided by a pre-trained model during generation. For example, landmark-based approaches [12, 44, 33, 38] detect predefined facial landmarks to transfer the facial pose and expression from a driving frame to the source image. Some other methods [22, 39, 35] use parameters from 3D face models [1, 8, 47] as motion descriptors to disentangle identity and pose. However, they normally cannot describe non-facial parts such as hair and neck, and their generation quality is limited by the pre-trained model performance. To address the issue, several methods that do not require any prior knowledge from pre-trained models are proposed. Monkey-Net [23] learns sparse motion-related keypoints in an unsupervised manner to describe object movements. FOMM [24] extends it with local affine transformation assumption around the keypoints to model complex motion. Subsequent works introduce more flexible mathematical models such as thin-plane spline transformation [43] and continuous piecewise-affine-based transformation [27] to increase motion estimation accuracy. Despite the expressiveness of the motion models, they cannot fully describe large head poses and delicate expression changes. MRFA [25] tackles the problem by building a correlation volume for each image pair and using it to refine the coarse motion flow iteratively. However, it only uses the warped image feature and a plain image generator for image generation, which may fail when facing extreme pose changes, as the appearance information from the one-shot source image is often insufficient. Some works [36, 21, 2] leverage the remarkable generative power of pre-trained StyleGAN [17, 18] for better image generation, but they usually have to balance the editability and fidelity. MCNet [15] learns a meta-memory bank of spatial facial features to compensate for the warped source features. Different from previous works, we compensate for both motion flows and warped source features with jointly learned multi-scale motion and appearance codebooks to boost the generation quality.
@@
Codebook Learning. Codebook learning aims to learn useful discrete representations with a fixed size. The learned codebook contains rich and compact information, which can facilitate various tasks such as image classification [3, 41], image synthesis [7, 4], blind face restoration [45, 10] and audio-driven talking head video generation [28]. Traditional methods often learn a codebook with unsupervised clustering such as k-means [5]. VQ-VAE [26] first incorporates vector quantization in a Variational Autoencoder to learn a codebook containing discrete representative input features. To build a context-rich codebook for images, VQGAN [7] further increases the compression rate and adds a discriminator and a perceptual loss on images reconstructed with the codes from the codebook. A transformer is later used to model the composition of the codes for high-resolution image synthesis. CodeFormer [45] uses a learned codebook of compressed high-quality face image features as discrete prior and predicts the code sequence based on the low-quality facial input for blind face restoration. LipFormer [28] learns two codebooks of the upper half face and the bottom half face respectively and predicts the lip codes from the input audio to generate a face video from the audio. We also adopt the idea of codebook learning, but we simultaneously learn multi-scale motion and appearance codebooks that store diverse motion and appearance patterns from the entire dataset during training to facilitate high-quality talking head video generation. The codebooks and the entire framework are trained together so that patterns useful for talking head video generation can be stored in and retrieved from the codebooks.
@@ 3 The Proposed Method
In this section, we will present the details of our framework. We jointly learn multi-scale motion and appearance codebooks with compensation for the motion and intermediate appearance features during generation. We adopt the Taylor expansion approximation method to learn the initial motion flow and the warping manner as same as Siarohin et al. [23] for video generation.
@@ 3.2.1 Multi-scale Code Allocation
We aim to refine the initial motion flow and the warped source features across all $N$ scales with multi-scale motion and appearance codebooks. As the scale increases, larger appearance features require finer motion flows for accurate warping and more detailed appearance information for appearance compensation. To provide sufficient motion and appearance compensation at larger scales, we introduce a code allocation scheme for multi-scale motion and appearance codebooks, which divides the codebook into multiple groups and allocates more codes for larger scales. We illustrate the scheme with the motion codebook in Fig. 5. The motion codebook $\mathcal{C}_{M}=\{\mathbf{m}_{k}\in\mathbb{R}^{d_{m}}\}_{k=1}^{K}$ contains $K$ codes, and we perform motion compensation at $N$ different scales. The $K$ codes are split into $N$ equal groups, each with $K/N$ codes. At scale $i$, the first $i$ groups, totaling $N_{m}^{i}=i\times K/N$ codes, are allocated for motion compensation. Similarly, for the appearance codebook $\mathcal{C}_{A}=\{\mathbf{a}_{k}\in\mathbb{R}^{d_{a}}\}_{k=1}^{T}$ containing $T$ codes, we also allocate the first $i$ groups, which are the first $N_{a}^{i}=i\times T/N$ codes for appearance compensation. This allows codes with smaller indices to capture general motion patterns or coarse appearance patterns shared across scales, while codes with larger indices to focus on finer motion patterns or facial details needed for larger scales. This maximizes the use of the two codebooks by sharing general information across scales while reserving space for scale-specific details.
@@ 3.2.2 Joint Codebook Learning
Motion Codebook Learning. At scale $i$, we use a CNN-based motion flow encoder $\mathcal{E}_{M}$ to map the input motion flow $\mathbf{M}^{i-1}\in\mathbb{R}^{h\times w\times 2}$ into a compact motion flow feature $\mathbf{F}_{m}^{i-1}\in\mathbb{R}^{h_{m}\times w_{m}\times d_{m}}$ where each unit is of length $d_{m}$ and captures a local motion flow pattern on $\mathbf{M}^{i-1}$. We quantize each of its spatial elements $\mathbf{F}_{m}^{i-1}(x,y)$ with the nearest code from $\mathcal{C}_{M}^{i}$ to obtain a quantized motion flow feature $\hat{\mathbf{F}}_{m}^{i-1}$:
@@
$\begin{split}\mathcal{L}_{vq,m}^{i}=&\lambda_{recon,m}||\hat{\mathbf{M}}^{i-1}-sg[\mathbf{M}^{i-1}]||_{1}\\ &+||sg[\mathcal{E}_{M}({\mathbf{M}}^{i-1})]-\hat{\mathbf{F}}_{m}^{i-1}||_{2}^{2}\\ &+\beta||sg[\hat{\mathbf{F}}_{m}^{i-1}]-\mathcal{E}_{M}(sg[\mathbf{M}^{i-1}])||_{2}^{2},\end{split}$ (3)
where $sg[\cdot]$ denotes the stop gradient operator, and $\lambda_{recon,m}$ and $\beta$ are the loss term weights. The first term represents the motion flow reconstruction loss, while the last two terms form a code-level loss [26] that minimizes the distance between the latent motion flow units and the motion codes. We stop the gradient of $\mathbf{M}^{i-1}$ to ensure that motion codebook learning at scale $i$ does not interfere with the training of the motion flow estimator or the compensated motion flows from prior scales. The overall loss function for motion codebook learning across all $N$ scales is $\mathcal{L}_{vq,m}=\sum_{i=1}^{N}\mathcal{L}_{vq,m}^{i}$.
@@ 3.2.3 Multi-scale Joint Codebook Compensation
Motion and Appearance Codebook Compensation Coupling. With the multi-scale motion and appearance codebooks, we couple motion and appearance codebook compensation at each scale to produce better image decoding features, as shown in Fig. 2. Specifically, at scale $i$, motion codebook compensation (MCC) refines the input motion flow $\mathbf{M}^{i-1}$ to obtain the compensated motion flow $\mathbf{M}^{i}$, which is used to warp the source feature $\mathbf{F}_{s}^{i}$. Appearance codebook compensation (ACC) then refines the warped source feature to produce the compensated appearance feature $\mathbf{F}_{c}^{i}$. If $i<N$, $\mathbf{M}^{i}$ is used as the input motion flow for the next scale. In this way, motion and appearance codebook compensation are coupled across scales, enabling more consistent motion and appearance for high-quality video generation.